When an ion‑selective electrode for X+ was immersed in 0.0652 M XCl, the measured potential was 0.0420 V . What is the concentration of X+ when the potential is 0.0570 V ? Assume that the electrode follows the Nernst equation, the temperature is at 25 °C, and that the activity coefficient of X+ is 1.
Answer #1

X+ +   e-   -----------> X

E = Eo + RT /n F ln [X+]

   = Eo + 8.314 x 298 / (1 x 96485) x ln [X+]

E = Eo + 0.02569 ln [X+]

when , [X+] = 0.0652 M , E = 0.0420 V

E = Eo + 0.02569 ln [X+]

0.0420 = Eo + 0.02569 ln [0.0652]

Eo = 0.112 V

when E = 0.0570 V

E = Eo + 0.02569 ln [X+]

0.0570 = 0.112 + 0.02569 ln [X+]

[X+] = 0.117 M